Steps performed in each sampling cycle 1. Place the pixel’s column on the address

bus.

2. Read the old value from DRAM and store it in a register.

3. Read the new value from the flash A2D.

4. Write the new value to the DRAM.

5. Find the difference between the old value and the new value, and pass the result on to the targeting unit.